Miss Takken heats a 300 g piece of iron to a temperature of 500 degrees Celsius and then plunges it into a vat of 30 L of water at 5 degrees Celsius. What is the final temperature of the system?



1
Expert's answer
2021-06-15T10:02:20-0400

The final temperature for water and iron will be the same, "t_i".


Heat given by iron to water:


"Q_1=c_im_i(t_{f}-t_{ii}),"

heat received by water:


"Q_2=c_wm_w(t_{f}-t_{iw}),"

and, as we know, thermal equilibrium is


"Q_1+Q_2=0,\\
c_im_i(t_{f}-t_{ii})+c_im_i(t_{f}-t_{iw})=0,\\\space\\
t_f=\frac{c_im_it_{ii}+c_wm_wt_{iw}}{c_im_i+c_wm_w}=5.5°\text C."